OpenStudy (anonymous):

0.9x+8≤7x-5 (simplify your answer; type an inequality symbol then type an integer or fraction

OpenStudy (anonymous):

7x-0.9x>=8+5 6.1x>=13 x>=13/6.1 use calculator to find the final answer...

OpenStudy (anonymous):

-6.1x ≤ -13 6.1x >= 13 x >= 130/61

OpenStudy (anonymous):

ok with that rachel

OpenStudy (anonymous):

let me check it.

OpenStudy (anonymous):

it says it is wrong.

OpenStudy (anonymous):

maybe decimal version?

OpenStudy (anonymous):

well 130 / 61 = 2/1311 correct to 4 dec places

OpenStudy (anonymous):

2.1311

OpenStudy (anonymous):

it says the correct answer is ≥16.25

OpenStudy (anonymous):

what???? i'll check that out

OpenStudy (anonymous):

k

OpenStudy (anonymous):

thats wrong rachel - try substituting 2.5 into the original expression we get 0.9(2.5) + 8 = 10.25 (LHS) and on RHS 7(2.5) - 5 = 12.5 10.5 < 12.5 - so that fits and if u plug in 2.311 they two sides are equal so our result is correct

OpenStudy (anonymous):

well thank you, just unfortunately its the schools website if they say it's wrong it's wrong. I do thakn you for your help though.

OpenStudy (anonymous):

yw - books websites, teachers - they all make mistakes sometimes - they certainly are wrong this time